# The 2D free particle in the phase space quantum mechanics

### v3

@@ Appendix B Calculation of the Moyal product of functions
$\text{sgn}(\tan\theta)\,\arccos|\cos\theta|=\theta-\pi\,\Big\lfloor\frac{\theta+\pi/2}{\pi}\Big\rfloor,$ (B.17)
where $\lfloor\cdot\rfloor$ denotes the floor function.
Changing the variable $\chi^{\prime\prime}\rightarrow\chi^{\prime\prime}-\chi$ we receive
@@
$\times\exp\bigg(-i(\widetilde{m}+\widetilde{m}^{\prime})\big(\pi\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\alpha+\pi/2}{\pi}\Big\rfloor\big)\bigg)$
$\times\exp\bigg(i(m+m^{\prime})\big(\pi\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\pi/2}{\pi}\Big\rfloor\big)\bigg)$
@@
It is necessary to discuss relations between the angles $\chi^{\prime\prime},\;\chi,\;\alpha$ which follow from the Heaviside functions. One can see that
$\alpha\in[0,\pi/2)$ $\Rightarrow(\chi^{\prime\prime}-\chi){\rm mod}\,2\pi\in[0,\pi/2-\alpha)\cup(3\pi/2,2\pi),$
$\alpha\in(\pi/2,\pi]$ $\Rightarrow(\chi^{\prime\prime}-\chi){\rm mod}\,2\pi\in(\pi/2,3\pi/2-\alpha),$
$\alpha\in(\pi,3\pi/2)$ $\Rightarrow(\chi^{\prime\prime}-\chi){\rm mod}\,2\pi\in(5\pi/2-\alpha,2\pi),$
$\alpha\in(3\pi/2,2\pi)$ $\Rightarrow(\chi^{\prime\prime}-\chi){\rm mod}\,2\pi\in[0,\pi/2)\cup(7\pi/2-\alpha,2\pi).$
In the cases first and fourth ones the values of the functions $\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\alpha+\pi/2}{\pi}\Big\rfloor$ and $\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\pi/2}{\pi}\Big\rfloor$ are both even numbers. In the cases second and third ones these values are odd.
Let us assume $m,\,m^{\prime},\,\widetilde{m},\,\widetilde{m}^{\prime}\in\mathbb{Z}$. Thus if $m^{\prime}=\widetilde{m}$, then
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}}{2\sqrt{H(E-H)}}e^{i(m-\widetilde{m}^{\prime})\chi}\delta(E-\widetilde{E})Y(E-H)$
$\times\sum_{\alpha\in\mathcal{A}}\exp\Big[-\frac{2i}{\hbar}L\tan(\alpha)\Big]e^{i(m+\widetilde{m}^{\prime})\alpha}|\pi-\alpha|(-1)^{(1-\text{sgn}(\cos\alpha))\,(m+\widetilde{m}^{\prime})/2}.$
In order to calculate the sum with respect to $\alpha\in\mathcal{A}$ we use the trigonometric identity $\tan\big(\arccos(x)\big)=\frac{\sqrt{1-x^{2}}}{x}.$ After some manipulations we finally obtain
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\pi N_{Emm^{\prime}}N_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}\frac{1}{\sqrt{H(E-H)}}\delta(E-\widetilde{E})Y(E-H)$
@@
$\times e^{i(m-\widetilde{m}^{\prime})\chi}\cos\left(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}-(m+\widetilde{m}^{\prime})\,\arccos\sqrt{\frac{H}{E}}\right)$
$=\frac{\pi N_{Emm^{\prime}}N_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}}{N_{Em\widetilde{m}^{\prime}}}\delta(E-\widetilde{E})W_{Em\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L).$ (B.19)
If we denote $W_{Em}:=W_{Emm}$, then for $m=m^{\prime}=\widetilde{m}^{\prime}$ Eq. (B.19) becomes
$(W_{Em}\star_{(\text{M})}W_{\widetilde{E}m}\big)(T,\,\chi,\,H,\,L)=\pi N_{Em}\delta(E-\widetilde{E})W_{Em}(T,\,\chi,\,H,\,L).$ (B.20)

### v4

@@ Acknowledgment
Conflict of interest statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data availability statement
No data were created or analyzed in this study.
@@ Appendix B Calculation of the Moyal product of functions
$\text{sgn}(\tan\theta)\,\arccos|\cos\theta|=\theta-\pi\,\Big\lfloor\frac{\theta+\frac{\pi}{2}}{\pi}\Big\rfloor,$ (B.17)
where $\lfloor\cdot\rfloor$ denotes the floor function. Hence and after changing the variable $\chi^{\prime\prime}\rightarrow\chi^{\prime\prime}-\chi$, we receive
@@
$\times\underbrace{\exp\bigg(-i(\widetilde{m}+\widetilde{m}^{\prime})\pi\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\alpha+\frac{\pi}{2}}{\pi}\Big\rfloor\bigg)\exp\bigg(i(m+m^{\prime})\big(\pi\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\frac{\pi}{2}}{\pi}\Big\rfloor\bigg)}_{=:\mathcal{C}_{\alpha}}$
@@
In order to discuss for which ranges of $\chi^{\prime\prime}-\chi$ the product $Y\bigg(\frac{\cos(\chi^{\prime\prime}-\chi+\alpha)}{\cos(\alpha)}\bigg)Y\bigg(\frac{\cos(\chi^{\prime\prime}-\chi)}{\cos(\alpha)}\bigg)$ equals $1$, we have to discuss four cases with respect to $\alpha$. We present the conclusions in table 1.
$\alpha$ $_{-} u+u_{+}$ $\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\alpha+\frac{\pi}{2}}{\pi}\Big\rfloor$ $\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\frac{\pi}{2}}{\pi}\Big\rfloor$ $\mathcal{C}_{\alpha}$
$\big(0,\,\frac{\pi}{2}\big)$ $\frac{\pi}{2}$ $\frac{\pi}{2}-\alpha$ 0 0 11
$\big(\frac{\pi}{2},\,\pi\big)$ $\frac{\pi}{2}$ $\frac{3\pi}{2}-\alpha$ 11 11 $e^{i\pi(-\widetilde{m}-\widetilde{m}^{\prime}+m+m^{\prime})}$
$\big(\pi,\,\frac{3\pi}{2}\big)$ $\frac{5\pi}{2}-\alpha$ $\frac{3\pi}{2}$ 11 33 $e^{i\pi(-3\widetilde{m}-3\widetilde{m}^{\prime}+m+m^{\prime})}$
$\big(\frac{3\pi}{2},\,2\pi\big)$ $\frac{3\pi}{2}-\alpha$ $\frac{\pi}{2}$ 0 22 $e^{i\pi(-2\widetilde{m}-2\widetilde{m}^{\prime})}$
Table 1: Four cases of range of $\alpha$ for which: if $(\chi^{\prime\prime}-\chi)\in(u_{-},\,u_{+})$, then $Y\bigg(\frac{\cos(\chi^{\prime\prime}-\chi+\alpha)}{\cos(\alpha)}\bigg)Y\bigg(\frac{\cos(\chi^{\prime\prime}-\chi)}{\cos(\alpha)}\bigg)=1$. Constant $\mathcal{C}_{\alpha}$ is defined in Eq. (B.18).
From Table 1, one concludes that we can factor exponents with $\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\alpha+\frac{\pi}{2}}{\pi}\Big\rfloor$ and $\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\frac{\pi}{2}}{\pi}\Big\rfloor$ out of the integral. Using indefinite integrals $\int\text{d}xe^{iax}=\frac{1}{a}e^{iax}$ if $a\neq 0$ and $\int\text{d}xe^{iax}=x$ if $a=0$ and information from the table, one can obtain that
$\times\sum_{\alpha\in\mathcal{A}}e^{i(m-m^{\prime}+\widetilde{m}+\widetilde{m}^{\prime})\alpha}\exp\Big[-\frac{2i}{\hbar}L\tan(\alpha)\Big]\mathcal{C}_{\alpha}\mathcal{F}_{\alpha},$ (B.19)
where for each $\alpha\in\mathcal{A}$, $A:=(\chi+u_{-})\text{mod}\,2\pi$ and $B:=(\chi+u_{+})\text{mod}\,2\pi$
$\mathcal{F}_{\alpha}:=\begin{cases}|\pi-\alpha|&,\,\widetilde{m}=m^{\prime},\\ \frac{1}{2i(\widetilde{m}-m^{\prime})}\big[e^{2i(\widetilde{m}-m^{\prime})B}-e^{2i(\widetilde{m}-m^{\prime})A}+Y(A-B)(e^{i4\pi(\widetilde{m}-m^{\prime})}-1)\big]&,\widetilde{m}\neq m^{\prime}.\end{cases}$ (B.20)
If we assume that $m,\,m^{\prime},\,\widetilde{m},\,\widetilde{m}^{\prime}\in\mathbb{Z}$ and use the trigonometric identity $\tan\big(\arccos(x)\big)=\frac{\sqrt{1-x^{2}}}{x}$, then Eq. (B.19) reduces to the form
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\pi N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\frac{1}{\sqrt{H(E-H)}}\delta(E-\widetilde{E})Y(E-H)$
@@
$\times e^{i(m-\widetilde{m}^{\prime})\chi}\cos\left(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}-(m+\widetilde{m}^{\prime})\,\arccos\sqrt{\frac{H}{E}}\right)\delta_{m^{\prime}\widetilde{m}}$
$=\frac{\pi N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{N_{Em\widetilde{m}^{\prime}}}\delta(E-\widetilde{E})W_{Em\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L)\delta_{m^{\prime}\widetilde{m}}$. (B.21)
In addition, if $m=m^{\prime}$ and $\widetilde{m}=\widetilde{m}^{\prime}$, then Eq. (B.21) becomes
$(W_{Em}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}}\big)(T,\,\chi,\,H,\,L)=\pi N_{Em}\delta(E-\widetilde{E})W_{Em}(T,\,\chi,\,H,\,L)\delta_{m\widetilde{m}}$. (B.22)
If at least one of the numbers $m,\,m^{\prime},\,\widetilde{m},\,\widetilde{m}^{\prime}$ belongs to $\mathbb{R}\setminus\mathbb{Z}$, then, in Eq. (B.19), we cannot simplify the coefficients $\mathcal{C}_{\alpha}$.